Solve the following equations show steps!!!
1. q(q+4)=0
2. 4k^2=5k

Answer by sachi(548)

1. q(q+4)=0
so either q=0
q+4=0 =>q=-4
so solutions are 0 or -4
2. 4k^2=5k
or 4k^2-5k=0
or k(4k-5)=0
or k=0
or 4k-5=0 =>k=5/4
so solutions are 0 or 5/4
ans